1. A chemist wishes to make 5.00g of CuCl(s) by the reaction:

CuCl2(aq) + Na2SO3(aq) + H2O(l) → CuCl(s) + Na2SO4(aq) + 2HCl(aq).How many grams of CuCl2 and of Na2CO3 should be used if the method gives an 85.0% yield, CuCl2 is to be the limiting reactant, and a 50.0% excess of Na2SO3 is to be used?




1
Expert's answer
2021-11-19T12:06:02-0500

Molar mass of CUCl2 = 134.45

50/100×134.45= 67.225g

Molar mass of Na2CO3= 105.99

= 105.99×50/100×0.85= 42.4g